Is this correct? Find three consecutive even integers such that the largest is three times the smallest.
3n+(n+2)+(n+4)

Answer by rapaljer(4671)

The largest is n+4, and the smallest is n. Therefore, the equation should be:
"Largest IS 3 times the smallest."
n+4 = 3(n)

Subtract n from each side:
n+4-n = 3n - n
4=2n

Divide both sides by 2:
2=n
The three consecutive even integers are 2, 4, and 6!

R^2 at SCC